# Beyond Baby Blues: The Child Penalty in Mental Health in Switzerland

### v1

@@ 3.3.1 Outcome Variables
The primary outcome variable is the mental health status of a woman, proxied by whether a woman has an antidepressant prescription (ATC Code: N06A) in a particular month. Since antidepressant prescriptions can last several months, the outcome variable is smoothed over the months the individual likely consumed the antidepressant. This smoothing is based on the number of prescribed packages and pills per package. Typically, patients consume one pill per day, so a package containing 30 pills is assumed to last one month. Additionally, if a woman lacks an antidepressant prescription for one or two months between two months with prescriptions, continuous consumption during these months is assumed. This smoothing approach provides a less volatile outcome variable without altering the results. [^15] A robustness check with the non-smoothed outcome can be found in Appendix LABEL:appendix_non_smoothed. Nevertheless, antidepressant prescriptions are potentially still a noisy proxy for mental health, and they most likely underestimate mental health problems, as they are only prescribed in severe cases.
In addition to the primary outcome, we consider other mental health proxies, such as the number of visits to a GP or psychiatrist. These measures are relevant since GPs or psychiatrists prescribe most antidepressants. It is essential to note that there are limitations to these outcomes. Increased GP visits may result from various factors unrelated to mental health, such as childcare-related illnesses, and time constraints, particularly in the first years after childbirth, might influence the number of visits to a psychiatrist and GP more than the antidepressant prescriptions. Furthermore, many GPs are located in a Health Maintenance Organisation (HMO) with other physicians, such as gynaecologists or specialists. We combine GP and HMO visits. Therefore, there is a slight measurement error in this outcome variable. Furthermore, the number of visits to the GP or HMO can only be identified consistently from 2012 to 2022 for the CSS data, which leads to a smaller sample size for this outcome. The outcome variables are measured relative to the event-time of birth $t$, with $t+1,t+2,\dots$ representing the months post-childbirth, and $t-1,t-2,\dots$ the months pre-childbirth.
@@
Note: The figure shows the percentage of women taking antidepressants around the time of birth ($t=0$: time of first birth). The red-shaded area illustrates the pregnancy, and the grey-shaded area represents the 95% confidence interval. ($N=25,120$)
To assess whether the increase in antidepressant use for women is due to an immediate rise after childbirth leading to drug dependence, we examine first-time antidepressant prescriptions. Since the prescription data only begins in 2012, identifying actual first-time prescriptions is challenging. As we move further in time, the likelihood increases that the observed prescriptions are genuinely first-time prescriptions, explaining the negative trend in antidepressant use among women without children (blue line) in Figure 2. To minimise the problem, we select women with at least one year without antidepressant prescriptions before their first recorded prescription. The number of women receiving their first antidepressant prescription stabilises at a consistent level in the years after childbirth. This suggests that the observed increase in antidepressant use is not only explainable by postpartum depression.
@@
Note: The figure shows the percentage of women taking for the first time antidepressants around the time of first birth ($t=0$). In the sample are only women who have at least one antidepressant prescription between 2012 and 2022 and who have had at least one year without antidepressant prescriptions before their first recorded prescription. The pink line shows the linear trend before and after childbirth for women giving birth in $t=0$. The blue line shows the linear trend for women who never have a child but get assigned random birthdates. The red-shaded area illustrates the pregnancy, and the grey-shaded area represents the 95% confidence interval. ($N=7,489$)
@@ 3.3.2 Covariates
We control for the individual’s age at their first birth in the analysis and use additional covariates to analyse heterogeneities. Controlling for age is important, as the risk of depression increases with age and because women who are mentally unhealthy might postpone having a child. Hence, women having a child at different ages may have different mental health trends. The average age of having a first child is around 29 years, and the age distribution is shown in Figure 5.181818The average age of Swiss women when they give birth for the first time is 30. However, the mean in our sample is lower due to selecting only women born between 1985 and 1995 and having the years from 2010 to 2022. By construction, the highest age at the birth of a first child is 37 years, and the lowest is 20 due to some sample restrictions.
@@
Accident coverage is a proxy for employment status, as it is typically needed by those unemployed, out of the labour force, or self-employed. At the same time, employers generally provide it for their employees. Figure 6 shows a notable increase in accident coverage following childbirth, suggesting a rise in non-employment among women. Around 80% of women do not have accident coverage before childbirth, implying an 80% employment rate. [^19] The decrease in accident coverage pre-birth is partly due to young women and men not yet having accident insurance because they are, for example, studying or doing an apprenticeship. This contrasts with the 90% pre-childbirth employment rate for women in Switzerland as reported by official statistics, indicating a discrepancy due to self-employment (BfS:2022). 8.1% of women without and 11.8% of women with children in Switzerland are self-employed. As depicted by Figure 6, accident coverage rises to about 30% (70% employed) post-childbirth, and further increases to 40%, suggesting 60% employment. Official figures show a decrease in employment to 80% after the first child and 69% after the second (BfS:2022). Thus, the proxy continuously underestimates employment by approximately ten percentage points (p.p.). This ten p.p. discrepancy is likely due to the exclusion of self-employed individuals. [^20] There is a concern that individuals might forget to cancel their accident insurance when they find a job, as it is inexpensive. However, the more common scenario is that people stop working after childbirth and need to sign up for accident coverage. Since accident insurance is mandatory and reminders are sent if it is not in place, this issue is less pronounced.
Figure 6: Descriptives: Accident coverage
@@
Note: The figure shows the proportion of women having accident coverage around the time of birth ($t=0$: time of first birth). The red shaded area illustrates the pregnancy, and the shaded area represents the 95% confidence interval. ($N=25,120$)
The left part of Table 3.3.2 shows the sociodemographic characteristics of women in the month they have a child, separated by the different samples used, whereas the right part of the table shows it separated by women having an antidepressant prescription or not.
@@
In the balanced sample, 19% of women receive subsidies for health care, 63% of women have more than one child, and 30% have a cesarean section. Furthermore, 78% of the women have German as a contact language, 19% have French, and 3% have Italian. The balanced and the unbalanced samples differ in some sociodemographic characteristics. For example, the balanced sample has a higher number of Swiss women and fewer women who speak English. A reason could be that some foreign women entered Switzerland later than 2012 and, hence, are not insured at the same insurance for the whole period. Similarly, some sociodemographic characteristics differ between the sample going until 2022 and the sample only going until 2019. On average, the latter has a lower age at first birth. By construction, the oldest woman giving birth can only be 34 years old instead of 37, which pushes the average age downwards. The unbalanced has more women with only one child by construction. We do not see all the years for the women in the unbalanced sample, hence, we miss some second births.
Comparing women who have an antidepressant prescription with those who do not, we see that slightly more women speaking French and fewer women speaking German have an antidepressant prescription. This pattern can also be seen in the regions of residence, as more women living in the French-speaking part of Switzerland have an antidepressant prescription compared to, for example, central Switzerland. As expected, women with an antidepressant prescription are more likely to have an accident insurance, meaning that they are more likely to not be employed. Furthermore, they are more likely to have a subsidy for health care, have a cesarean section and only have one child.

### v2

@@ 3.3.1 Outcome Variables
The primary outcome variable is the mental health status of a woman, proxied by whether a woman has an antidepressant prescription (ATC Code: N06A) in a particular month. Since antidepressant prescriptions can last several months, the outcome variable is smoothed over the months the individual likely consumed the antidepressant. This smoothing is based on the number of prescribed packages and pills per package. Typically, patients consume one pill per day, so a package containing 30 pills is assumed to last one month. Additionally, if a woman lacks an antidepressant prescription for one or two months between two months with prescriptions, continuous consumption during these months is assumed. This smoothing approach provides a less volatile outcome variable without altering the results. [^16] A robustness check with the non-smoothed outcome can be found in Appendix LABEL:appendix_non_smoothed. Nevertheless, antidepressant prescriptions are potentially still a noisy measure of mental health, and they most likely underestimate mental health problems, as not everyone with mental health issues receives medication. Antidepressants are typically prescribed for moderate to severe cases, but many individuals with milder symptoms rely on therapy, other treatments, or do not seek treatment at all (Obsan:2022).
In addition to the primary outcome, we consider other mental health proxies, such as the number of visits to a GP or psychiatrist. These measures are relevant since GPs or psychiatrists prescribe most antidepressants. It is essential to note that there are limitations to these outcomes. Increased GP visits may result from various factors unrelated to mental health, such as childcare-related illnesses, and time constraints, particularly in the first years after childbirth, might influence the number of visits to a psychiatrist and GP more than the antidepressant prescriptions. Furthermore, many GPs are located in a Health Maintenance Organisation (HMO) with other physicians, such as gynaecologists or specialists. We combine GP and HMO visits. Therefore, there is a slight measurement error in this outcome variable. Moreover, the number of visits to the GP or HMO can only be identified consistently from 2012 to 2022 for the CSS data, which leads to a smaller sample size for this outcome. The outcome variables are measured relative to the event-time of birth $t$, with $t+1,t+2,\dots$ representing the months post-childbirth, and $t-1,t-2,\dots$ the months pre-childbirth.
@@
Note: The figure shows the percentage of women taking antidepressants around the time of birth ($t=0$: time of first birth). The red-shaded area illustrates the pregnancy, and the grey-shaded area represents the 95% confidence interval. The slow decrease in antidepressants at the start of the pregnancy and also the slow increase after giving birth are partly coming from the smoothing of the antidepressant prescriptions. Compared to Denmark or Austria, the antidepressant prescription rate is lower before childbirth. A reason, therefore, could be that even after monthly health insurance payments, Swiss individuals have to pay for medical visits and drugs up to a yearly deductible (often CHF 2,500). ($N=25,120$)
To assess whether the increase in antidepressant use for women is due to an immediate rise after childbirth leading to drug dependence, we examine first-time antidepressant prescriptions. Since the prescription data only begins in 2012, identifying actual first-time prescriptions is challenging. As we move further in time, the likelihood increases that the observed prescriptions are genuinely first-time prescriptions, explaining the negative trend in antidepressant use among women without children (blue line) in Figure 2.181818For the women without children placebo dates for the first birth are randomly drawn. We follow Ahammer:2023 and approximate the factual distribution of age at first birth by a birth-cohort specific log-normal distribution $LN(\mu,\sigma^{2})$. The month is randomly drawn from a uniform distribution. To minimise this issue, we select women who had at least one year without antidepressant prescriptions before their first recorded prescription. We find that the number of women receiving their first antidepressant prescription stabilizes at a consistent level in the years following childbirth. This suggests that the observed increase in antidepressant use is not solely driven by postpartum depression but persists even two to six years after childbirth, indicating a longer-term effect.
@@
Note: The figure shows the percentage of women taking for the first time antidepressants around the time of first birth ($t=0$). In the sample are only women who have at least one antidepressant prescription between 2012 and 2022 and who have had at least one year without antidepressant prescriptions before their first recorded prescription. The pink line shows the linear trend before and after childbirth for women giving birth in $t=0$. The blue line shows the linear trend for women who never have a child but get assigned random birthdates. The factual distribution of age at first birth for women without children has been approximated by a birth-cohort specific log-normal distribution $LN(\mu,\sigma^{2})$. The month is randomly drawn from a uniform distribution. The red-shaded area illustrates the pregnancy, and the grey-shaded area represents the 95% confidence interval. ($N=7,489$)
@@ 3.3.2 Covariates
We control for the individual’s age at their first birth in the analysis and use additional covariates to analyse heterogeneities. Controlling for age is important, as the risk of depression increases with age and because women who are mentally unhealthy might postpone having a child. Consequently, women giving birth at different ages may exhibit different underlying trends in mental health, which could bias estimates if not accounted for. The average age of having a first child is around 29 years, and the age distribution is shown in Figure 5.202020The average age of Swiss women when they give birth for the first time is 30. However, the mean in our sample is lower due to selecting only women born between 1985 and 1995 and having the years from 2010 to 2022. By construction, the highest age at the birth of a first child is 37 years, and the lowest is 20 due to some sample restrictions.
@@
Employment status can be identified by looking at the accident coverage of the health insurance. Employers generally provide it for their employees, hence, employed individuals do not have accident coverage in the health insurance. On the other hand, it is needed by those unemployed, out of the labour force or self-employed. Figure 6 shows a notable decrease in employement following childbirth. Around 80% of women are employed before childbirth. [^21] The increase in employment pre-birth is partly due to young women and men not yet started to work as they still are, for example, studying or doing an apprenticeship. This contrasts with the 90% pre-childbirth employment rate for women in Switzerland as reported by official statistics, indicating a discrepancy due to self-employment (BfS:2022). 8.1% of women without and 11.8% of women with children in Switzerland are self-employed. As depicted by Figure 6, employment rates drop to about 70% post-childbirth, and further reduce to 60%. Official figures show a decrease in employment to 80% after the first child and 69% after the second (BfS:2022). Thus, the proxy continuously underestimates employment by approximately ten percentage points (p.p.). This ten p.p. discrepancy is likely due to the exclusion of self-employed individuals. [^22] There is a concern that individuals might forget to cancel their accident insurance when they find a job, as it is inexpensive. However, the more common scenario is that people stop working after childbirth and need to sign up for accident coverage. Since accident insurance is mandatory and reminders are sent if it is not in place, this issue is less pronounced.
Figure 6: Descriptives: Employment
@@
Note: The figure shows the proportion of women employed around the time of birth ($t=0$: time of first birth). The red shaded area illustrates the pregnancy, and the shaded area represents the 95% confidence interval. ($N=25,120$)
The left part of Table 3.3.2 presents the sociodemographic characteristics of women in the month they give birth, across the different samples used in the analysis. This helps to evaluate whether the sample selection might introduce bias and whether robustness checks are warranted. The right part of the table compares women with and without an antidepressant prescription prior to childbirth, highlighting potentially meaningful differences between these groups that motivate further investigations of heterogeneous treatment effects.
@@
In the balanced sample, 19% of women receive subsidies for health care, 63% of women have more than one child, and 30% have a cesarean section. Furthermore, 78% of the women have German as a contact language, 19% have French, and 3% have Italian. The balanced and the unbalanced samples differ in some sociodemographic characteristics. For example, the balanced sample has a higher number of Swiss women and fewer women who speak English. A reason could be that some foreign women entered Switzerland later than 2012 and, hence, are not insured at the same insurance for the whole period. Similarly, some sociodemographic characteristics differ between the sample going until 2022 and the sample only going until 2019. On average, the latter has a lower age at first birth. By construction, the oldest woman giving birth can only be 34 years old instead of 37, which pushes the average age downwards. The unbalanced sample has more women with only one child by construction. We do not see all the years for the women in the unbalanced sample, hence, we miss some second births. As there are some differences between the two samples, the unbalanced sample is used as a robustness check (see Section LABEL:robustness:unbalanced_sample).